OpenStudy (anonymous):

simplify 1/x + 2/9 = 1/3

OpenStudy (anonymous):

\[\frac{1}{x}+\frac{2}{9}=\frac{1}{3}\]\[\frac{1}{x}=\frac{1}{3}-\frac{2}{9}\] \[\frac{1}{x}=\frac{3}{9}-\frac{2}{9}\]\[\frac{1}{x}=\frac{1}{9}\]

OpenStudy (anonymous):

Cross multiply to get the answer

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

so its 9/x

OpenStudy (anonymous):

x=9